40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Aug 7 14:11:53 2011, Jenne, Update, PEM, Particle counts through the roof ParticleCount_High_7Aug2011.png
    Reply  Mon Aug 8 14:21:03 2011, steve, Update, PEM, Particle counts controlled P1080149.JPGP1080150.JPG
Message ID: 5134     Entry time: Sun Aug 7 14:11:53 2011     Reply to this: 5140
Author: Jenne 
Type: Update 
Category: PEM 
Subject: Particle counts through the roof 

[Jenne, Kiwamu]

While Kiwamu was finalizing the X green alignment, I started to prepare to remove the ETMY door, and begin checking out its OSEMs, etc, so we could start moving it to it's new place, and figure out why it's been wonky for a while.  I ran the particle counter, and we have a factor of ~5 more particles than normal.  Kiwamu and I agreed not to open ETMY.  Since we had briefly opened the IOO and Output Optics chambers to check the X green's position on the PSL table, we immediately shut those doors.  They were probably open for ~15 minutes or so.  (Yes Steve, we should have checked before opening any doors, but at least we remembered to check at all, and the doors were only open for a few minutes rather than for a few hours.)

I attach a 24hrs trend of the particle counts, for reference.  It looks like it's been a little high for a while, but today it's really dirty in the air.

Attachment 1: ParticleCount_High_7Aug2011.png  11 kB  | Hide | Hide all
ParticleCount_High_7Aug2011.png
ELOG V3.1.3-